OpenStudy (anonymous):

find f(x) and g(x) so the function can be expressed as y=f(g(x)) y=(7/x^2) +10

OpenStudy (anonymous):

g(x)=7/x^2 f(x)=g(x)+10

OpenStudy (anonymous):

do you understand @psila

OpenStudy (anonymous):

do you mind explaining it? i'm sorry!

OpenStudy (anonymous):

so its a question you have to look at closely and think. y=f(g(x)) so you see that that g(x) is inside the f(x) function, so the whole thing is the f(x) thus making the part in parenthesis the g(x) function

OpenStudy (anonymous):

i hope that makes sense @psila

OpenStudy (anonymous):

it does! thank you so much!

OpenStudy (anonymous):

no problem :)